# Unified Primitive Proxies for Structured Shape Completion

### v1

@@ 1 Introduction
Occlusions and limited sensor coverage often leave 3D scans incomplete. Completing the missing geometry allows robots to plan stable grasps, enables autonomous vehicles to perceive hidden traffic, and supports digitizing heritage artifacts without repeated acquisitions [tabib2023defi, tsesmelis2024re, zheng2024towards, varley2017shape].
Despite advances, most shape completion methods still optimize pointwise discrepancies [yuan2018pcn, yu2021pointr, yu2023adapointr] or their variants [wu2021density, lin2023hyperbolic, lin2023infocd]. These objectives capture local geometry but convey little about the structural regularities required by many downstream tasks [botsch2010polygon, koch2019abc]. In contrast, primitive assembly models the surface as a compact, topologically consistent collection of parametric primitives for structured, interpretable geometry [nan2017polyfit, bauchet2020kinetic, sulzer2024concise, jiang2023primfit]. However, the common recipe of completing first and assembling later is underconstrained, because assembly solvers expect structured input that pointwise completion does not provide. It is therefore preferable to predict structure jointly with completion.
Toward structured shape completion that directly supports primitive assembly, a straightforward approach is a two-stage cascade that first regresses primitive parameters and then enforces inlier points, typically limited to plane-only primitives [chen2025parametric]. In practice, this rigid formulation tends to overfit well-supported regions and degrade when evidence is sparse. It can also propagate early errors in primitive count or parameters into the association step, which weakens later supervision. These limitations motivate a formulation in which point completion and primitive inference are optimized in a more coordinated way.
@@ 2 Related Work
Early approaches to shape completion used volumetric CNNs, but voxel representations suffer from discretization artifacts and high memory costs at fine resolutions [wu2015, dai2017, xie2020grnet, han2017high]. The introduction of PointNet enabled direct processing of unordered point sets [qi2017pointnet], which led to a broad family of point-based completion networks [qi2017pointnet++, wang2019dynamic, thomas2019kpconv]. Most recent methods, including PoinTr [yu2021pointr], AdaPoinTr [yu2023adapointr], ODGNet [cai2024orthogonal], SymmComplete [yan2025symmcompletion], and others [yuan2018pcn, yang2018foldingnet, xiang2021snowflakenet, zhao2022seedformer, tang2022lake, yan2022fbnet, chen2023anchorformer, zhu2023csdn, tesema2023point, lee2024proxyformer], still minimize pointwise discrepancies or their variants [wu2021density, lin2023hyperbolic, lin2023infocd], so they recover local geometry but leave higher-level structure underconstrained. PaCo moves toward structured shape completion by first predicting plane parameters and then enforcing inlier membership in a cascade, yet this design remains sensitive to sparse evidence and early errors, and is restricted to plane-only primitives [chen2025parametric]. In contrast, our approach unifies shape completion and structural reasoning in a single network with coordinated pathways, enabling consistent optimization across multiple primitive families.
@@
Unlike generic surface reconstruction that targets dense meshes [kazhdan2013screened, erler2020points2surf, huang2022neural, huang2024surface], primitive assembly reconstructs surfaces by arranging primitives under geometric and topological constraints. Although recent neural approaches explore learned decompositions and constructive modeling [tulsiani2017learning, chen2020bspnet, ren2021csg, yu2022capri, li2023secad, liu2024point2cad], primitive assembly remains the practical standard when topology control, editability, and compatibility are required [botsch2010polygon, kaiser2019survey, chen2024polygnn, chen2022points2poly, chen2025abspy]. Representative solvers include PolyFit [nan2017polyfit], KSR [bauchet2020kinetic], and COMPOD [sulzer2024concise] for polygonal surfaces, and PrimFit [jiang2023primfit] for more general primitive families. However, their performance degrades under partial observations because assembly depends on reliable, complete primitives, which motivates learning structured representations from partial scans that supply assembly-ready primitives.
@@
Primitive extraction is often cast as instance segmentation. Traditional pipelines rely on geometric model fitting and constraints, but lack learned inductive biases [schnabel2007efficient, li2011globfit, yu2022finding]. Learning-based methods later adopted a clustering paradigm, where a shared backbone first produces pointwise features and a subsequent stage groups points into primitive instances [li2019supervised, sharma2020parsenet, le2021cpfn, yan2021hpnet]. This paradigm led benchmarks for several years [jiang2020pointgroup, lahoud20193d, chen2021hierarchical, liang2021instance, vu2022softgroup]. More recent Transformer-based models predict masks directly with instance queries and avoid hand-crafted grouping [schult2022mask3d, sun2023superpoint, lai2023mask, lu2023query]. However, they rarely encode primitive-specific priors and typically assume complete, fixed point sets, which hinders transfer to completion settings with inconsistent targets. In contrast, we infer primitives in tandem with completion, jointly predicting memberships from incomplete scans while reasoning about missing geometry.
@@ 3.1 Coordinated Pathways
We instantiate the point pathway $f_{\mathrm{point}}$ with AdaPoinTr [yu2023adapointr], though in principle any pointwise shape completion network could be used. This pathway recovers a local point patch from each feature $\mathbf{t}_{u}$ and aggregates them into the completed points $\mathcal{\hat{Y}}\subset\mathbb{R}^{3}$:
@@
where $\mathbf{y}_{j}^{u}$ denotes a point in the patch decoded from $\mathbf{t}^{u}$.
@@ 3.3 Optimization
Here, $\alpha_{1},\alpha_{2},\alpha_{3}$ and $\lambda$ are balancing weights. The semantic term encourages correct primitive type prediction. The membership term enforces point-primitive membership consistency using cross-entropy and Dice losses [deng2018learning]. The geometry term aligns predicted inliers with the ground truth in both Chamfer distance and parameters. The optimal bipartite matching $\mathcal{M}$ is obtained with the Hungarian algorithm [kuhn1955hungarian]. The overall loss combines the matched primitive costs with an object-level distance from the point pathway:
@@
At inference time, inspired by practices in instance segmentation [cheng2021per, schult2022mask3d], we score each predicted primitive by combining its semantic confidence with the reliability of its inliers:
@@
Primitives with $s_{k}\geq 0.5$ are retained and passed to the downstream assembly solver. Architectural and implementation details are provided in the Appendix.
Figure 3: Comparison to completion baselines on ABC-multi. For each baseline, completed points are paired with its best-performing primitive extractor (  RANSAC [schnabel2007efficient],  HPNet [yan2021hpnet],  PTv3 [wu2024ptv3]). UniCo recovers extractor-free, assembly-ready primitive structures.
@@ 4.1 Setup and Protocol
We evaluate on three datasets for a comprehensive assessment. ABC-multi is a curated subset of 30,000 watertight CAD models from the ABC dataset [koch2019abc], with 5,000 reserved for evaluation. It spans planes, cylinders, spheres, and cones, stressing completion with mixed primitive types. ABC-plane [chen2025parametric] is an existing plane-only dataset with over 15,000 CAD models. We use it to assess behavior in plane-dominant settings and to compare with polygonal pipelines. For real-world evaluation, we use Building-PCC [gao2024building], an airborne LiDAR dataset of about 50,000 urban buildings with realistic noise and occlusions. Following prior works [yu2023adapointr, chen2025parametric], ABC-multi and ABC-plane use 2,048 input points down-sampled from partial points at 25%, 50%, and 75% incompleteness, and 8,192 target points per shape, while Building-PCC uses native point counts.
@@
We compare three groups of baselines. For completion, we evaluate GRNet [xie2020grnet], PoinTr [yu2021pointr], AdaPoinTr [yu2023adapointr], ODGNet [cai2024orthogonal], SymmComplete [yan2025symmcompletion], and PaCo [chen2025parametric]. Unless stated otherwise, all methods are trained to convergence. For primitive extraction, on ABC-multi we fit primitives to completed points using RANSAC [schnabel2007efficient], HPNet [yan2021hpnet], and PTv3 [wu2024ptv3], followed by PrimFit [jiang2023primfit] for assembly. On ABC-plane and Building-PCC, we use GoCoPP [yu2022finding] to extract primitives and then apply PolyFit [nan2017polyfit], KSR [bauchet2020kinetic], and COMPOD [sulzer2024concise] for assembly. For reconstruction, we include BSP-Net [chen2020bspnet] and Point2CAD [liu2024point2cad], evaluated on both raw and completed points. For multi-stage pipelines (e.g., completion followed by reconstruction), we compose best models at each stage to obtain competitive baselines.
@@ 4.2 Mixed-Type Primitive Results
Table 1: Comparison with completion baselines on ABC-multi. Best scores are bold, second best are underlined. Baseline methods require a separate primitive extractor, whereas UniCo produces structured completion directly. All results are reconstructed with the same PrimFit solver [jiang2023primfit]. CD, HD, and FR values are scaled by 100.
Method Pointwise RANSAC [schnabel2007efficient] HPNet [yan2021hpnet] PTv3 [wu2024ptv3]
@@
GRNet [xie2020grnet] 1.177 0.565     18.61 24.71 0.543 100.00     18.61 24.71 0.543 100.00     18.61 24.71 0.543 100.00
PoinTr [yu2021pointr] 0.719 0.764     6.58 22.62 0.814 17.07     7.56 21.38 0.779 26.70     6.58 19.41 0.791 11.27
AdaPoinTr [yu2023adapointr] 0.625 0.824     6.81 21.24 0.815 19.77     4.41 13.36 0.872 8.97     5.58 16.80 0.821 9.83
ODGNet [cai2024orthogonal] 0.632 0.850     4.80 22.15 0.868 0.39     4.33 13.63 0.873 7.41     5.48 16.79 0.823 9.18
SymmComplete [yan2025symmcompletion] 0.487 0.825     7.54 21.42 0.796 26.13     4.57 13.58 0.865 9.84     5.93 17.58 0.812 11.50
@@
PrimFit [jiang2023primfit] R 7.80 25.17 0.788 15.96
Point2CAD [liu2024point2cad] R 8.69 30.66 0.779 15.91
BSP-Net [chen2020bspnet] R 8.35 16.73 0.740 20.52
@@
Figure 5: Robustness to missing data, transforms, and noise. Left: as incompleteness grows from 25% to 75%, UniCo maintains lower CD and higher NC than pointwise baselines. Middle: under the debiased normalization protocol [wang2025simeco], UniCo remains stable with low pose and scale bias compared to baselines. Right: under Gaussian jitter of 1–3%, performance degrades gracefully.
@@
We stress-test UniCo under increasing incompleteness, a debiased normalization, and Gaussian jitter, as presented in Fig. 5. All networks are trained for 200 epochs. As incompleteness rises from 25% to 75%, UniCo’s CD increases only from 1.8 to 2.7, whereas strong pointwise baselines double their error to about 6.0. NC for UniCo drops from 0.95 to 0.91, whereas the baselines drop to 0.82. Under a debiased normalization that removes pose and scale canonicalization [wang2025simeco], UniCo degrades to CD 3.9 with NC 0.88, whereas the baselines deteriorate to CD above 14 with NC below 0.65, indicating that UniCo carries substantially less pose and scale bias. With Gaussian noise up to 3%, UniCo remains stable. At a heavy 5% noise level, CD and NC degrade to 3.2 and 0.88, respectively. Fig. 6 further shows view-to-view consistency, where different partial inputs yield a consistent primitive set that supports reliable assembly.
@@
To better understand what drives the strong assembly performance, we further evaluate primitive quality following the established protocol [guo2022complexgen, liu2024point2cad, li2019supervised]. For all competing methods, primitives are extracted with HPNet [yan2021hpnet] for a feasible and fair comparison. Predicted primitives are then matched to ground truth via Hungarian matching. As shown in Tab. 3, we report F-Score@1% (F1), type accuracy (Type), axis difference (Axis), residual error (Res), and coverage (Cov). Axis measures the plane normal or the symmetry axis of cylinders and cones. Res evaluates the fit on 512 points sampled from the ground truth primitive. Cov is the fraction of these points that lie within a distance of 0.01 from the predicted primitive. UniCo attains the best scores on all primitive metrics, consistent with its stronger reconstruction performance.
@@
GRNet [xie2020grnet] 0.215 29.65 25.97 7.12 24.41
PoinTr [yu2021pointr] 0.509 64.41 18.24 2.86 61.60
AdaPoinTr [yu2023adapointr] 0.643 79.79 11.34 1.48 78.99
ODGNet [cai2024orthogonal] 0.659 75.52 12.24 1.78 75.85
SymmComplete [yan2025symmcompletion] 0.629 78.48 12.72 1.54 77.60
@@
Method Year PolyFit [nan2017polyfit] KSR [bauchet2020kinetic] COMPOD [sulzer2024concise]
@@
GRNet [xie2020grnet] 2020 11.98 19.84 0.769 29.61 9.18 22.01 0.822 10.82 14.19 25.72 0.738 13.15
PoinTr [yu2021pointr] 2021 10.57 16.43 0.822 25.92 8.14 16.33 0.780 30.90 7.82 16.44 0.774 31.34
AdaPoinTr [yu2023adapointr] 2023 3.16 7.36 0.920 5.89 3.24 8.86 0.927 0.27 3.25 8.84 0.921 1.32
ODGNet [cai2024orthogonal] 2024 2.73 6.41 0.933 4.28 2.90 8.56 0.934 0.36 3.22 8.01 0.927 1.05
SymmComplete [yan2025symmcompletion] 2025 3.21 9.15 0.920 4.27 4.23 14.21 0.907 0.38 4.56 14.02 0.895 4.14
PaCo [chen2025parametric] 2025 1.87 4.09 0.943 0.48 1.91 4.14 0.940 0.25 1.94 4.42 0.940 0.25

### v2

@@ 1 Introduction
Occlusions and limited sensor coverage often leave 3D scans incomplete. Completing the missing geometry allows robots to plan stable grasps, enables autonomous vehicles to perceive hidden traffic, and supports digitizing heritage artifacts without repeated acquisitions [47, 51, 73, 53].
Despite advances, most shape completion methods still optimize pointwise discrepancies [71, 69, 70] or their variants [57, 34, 35]. These objectives capture local geometry but convey little about the structural regularities required by many downstream tasks [2, 24]. In contrast, primitive assembly models the surface as a compact, topologically consistent collection of parametric primitives for structured, interpretable geometry [38, 1, 45, 20]. However, the common recipe of completing first and assembling later is underconstrained, because assembly solvers expect structured input that pointwise completion does not provide. It is therefore preferable to predict structure jointly with completion.
Toward structured shape completion that directly supports primitive assembly, a straightforward approach is a two-stage cascade that first regresses primitive parameters and then enforces inlier points, typically limited to plane-only primitives [7]. In practice, this rigid formulation tends to overfit well-supported regions and degrade when evidence is sparse. It can also propagate early errors in primitive count or parameters into the association step, which weakens later supervision. These limitations motivate a formulation in which point completion and primitive inference are optimized in a more coordinated way.
@@ 2 Related Work
Early approaches to shape completion used volumetric CNNs, but voxel representations suffer from discretization artifacts and high memory costs at fine resolutions [60, 12, 62, 17]. The introduction of PointNet enabled direct processing of unordered point sets [39], which led to a broad family of point-based completion networks [40, 55, 50]. Most recent methods, including PoinTr [69], AdaPoinTr [70], ODGNet [3], SymmComplete [63], and others [71, 66, 61, 72, 48, 65, 9, 74, 49, 29], still minimize pointwise discrepancies or their variants [57, 34, 35], so they recover local geometry but leave higher-level structure underconstrained. PaCo moves toward structured shape completion by first predicting plane parameters and then enforcing inlier membership in a cascade, yet this design remains sensitive to sparse evidence and early errors, and is restricted to plane-only primitives [7]. In contrast, our approach unifies shape completion and structural reasoning in a single network with coordinated pathways, enabling consistent optimization across multiple primitive families.
@@
Unlike generic surface reconstruction that targets dense meshes [23, 14, 18, 19], primitive assembly reconstructs surfaces by arranging primitives under geometric and topological constraints. Although recent neural approaches explore learned decompositions and constructive modeling [52, 10, 41, 67, 31, 36], primitive assembly remains the practical standard when topology control, editability, and compatibility are required [2, 22, 6, 5, 8]. Representative solvers include PolyFit [38], KSR [1], and COMPOD [45] for polygonal surfaces, and PrimFit [20] for more general primitive families. However, their performance degrades under partial observations because assembly depends on reliable, complete primitives, which motivates learning structured representations from partial scans that supply assembly-ready primitives.
@@
Primitive extraction is often cast as instance segmentation. Traditional pipelines rely on geometric model fitting and constraints, but lack learned inductive biases [42, 32, 68]. Learning-based methods later adopted a clustering paradigm, where a shared backbone first produces pointwise features and a subsequent stage groups points into primitive instances [30, 44, 28, 64]. This paradigm led benchmarks for several years [21, 26, 4, 33, 54]. More recent Transformer-based models predict masks directly with instance queries and avoid hand-crafted grouping [43, 46, 27, 37]. However, they rarely encode primitive-specific priors and typically assume complete, fixed point sets, which hinders transfer to completion settings with inconsistent targets. In contrast, we infer primitives in tandem with completion, jointly predicting memberships from incomplete scans while reasoning about missing geometry.
@@ 3.1 Coordinated Pathways
We instantiate the point pathway $f_{\mathrm{point}}$ with AdaPoinTr [70], though in principle any pointwise shape completion network could be used. This pathway recovers a local point patch from each feature $\mathbf{t}^{u}$ and aggregates them into the completed points $\mathcal{\hat{Y}}\subset\mathbb{R}^{3}$:
@@
where $\hat{\mathbf{y}}_{j}^{u}$ denotes a point in the patch decoded from $\mathbf{t}^{u}$.
@@ 3.3 Optimization
Here, $\alpha_{1},\alpha_{2},\alpha_{3}$ and $\lambda$ are balancing weights. The semantic term encourages correct primitive type prediction. The membership term enforces point-primitive membership consistency using cross-entropy and Dice losses [13]. The geometry term aligns predicted inliers with the ground truth in both Chamfer distance and parameters. The optimal bipartite matching $\mathcal{M}$ is obtained with the Hungarian algorithm [25]. The overall loss combines the matched primitive costs with an object-level distance from the point pathway:
@@
At inference time, inspired by practices in instance segmentation [11, 43], we score each predicted primitive by combining its semantic confidence with the reliability of its inliers:
@@
Primitives with $s_{k}>0.5$ are retained and passed to the downstream assembly solver. Architectural and implementation details are provided in the Appendix.
Figure 3: Comparison to completion baselines on ABC-multi. For each baseline, completed points are paired with its best-performing primitive extractor (  RANSAC [42],  HPNet [64],  PTv3 [59]). UniCo recovers extractor-free, assembly-ready primitive structures.
@@ 4.1 Setup and Protocol
We evaluate on three datasets for a comprehensive assessment. ABC-multi is a curated subset of 30,000 watertight CAD models from the ABC dataset [24], with 5,000 reserved for evaluation. It spans planes, cylinders, spheres, and cones, stressing completion with mixed primitive types. ABC-plane [7] is an existing plane-only dataset with over 15,000 CAD models. We use it to assess behavior in plane-dominant settings and to compare with polygonal pipelines. For real-world evaluation, we use Building-PCC [15], an airborne LiDAR dataset of about 50,000 urban buildings with realistic noise and occlusions. Following prior works [70, 7], ABC-multi and ABC-plane use 2,048 input points down-sampled from partial points at 25%, 50%, and 75% incompleteness, and 8,192 target points per shape, while Building-PCC uses native point counts.
@@
We compare three groups of baselines. For completion, we evaluate GRNet [62], PoinTr [69], AdaPoinTr [70], ODGNet [3], SymmComplete [63], and PaCo [7]. Unless stated otherwise, all methods are trained to convergence. For primitive extraction, on ABC-multi we fit primitives to completed points using RANSAC [42], HPNet [64], and PTv3 [59], followed by PrimFit [20] for assembly. On ABC-plane and Building-PCC, we use GoCoPP [68] to extract primitives and then apply PolyFit [38], KSR [1], and COMPOD [45] for assembly. For reconstruction, we include BSP-Net [10] and Point2CAD [36], evaluated on both raw and completed points. For multi-stage pipelines (e.g., completion followed by reconstruction), we compose best models at each stage to obtain competitive baselines.
@@ 4.2 Mixed-Type Primitive Results
Table 1: Comparison with completion baselines on ABC-multi. Best scores are bold, second best are underlined. Baseline methods require a separate primitive extractor, whereas UniCo produces structured completion directly. All results are reconstructed with the same PrimFit solver [20]. CD, HD, and FR values are scaled by 100.
Method Pointwise RANSAC [42] HPNet [64] PTv3 [59]
@@
GRNet [62] 1.177 0.565     18.61 24.71 0.543 100.00     18.61 24.71 0.543 100.00     18.61 24.71 0.543 100.00
PoinTr [69] 0.719 0.764     6.58 22.62 0.814 17.07     7.56 21.38 0.779 26.70     6.58 19.41 0.791 11.27
AdaPoinTr [70] 0.625 0.824     6.81 21.24 0.815 19.77     4.41 13.36 0.872 8.97     5.58 16.80 0.821 9.83
ODGNet [3] 0.632 0.850     4.80 22.15 0.868 0.39     4.33 13.63 0.873 7.41     5.48 16.79 0.823 9.18
SymmComplete [63] 0.487 0.825     7.54 21.42 0.796 26.13     4.57 13.58 0.865 9.84     5.93 17.58 0.812 11.50
@@
PrimFit [20] R 7.80 25.17 0.788 15.96
Point2CAD [36] R 8.69 30.66 0.779 15.91
BSP-Net [10] R 8.35 16.73 0.740 20.52
CARPRI-Net [67] R 7.59 15.18 0.770 19.27
@@
Figure 5: Robustness to missing data, transforms, and noise. Left: as incompleteness grows from 25% to 75%, UniCo maintains lower CD and higher NC than pointwise baselines. Middle: under the debiased normalization protocol [56], UniCo remains stable with low pose and scale bias compared to baselines. Right: under Gaussian jitter of 1–3%, performance degrades gracefully.
@@
We stress-test UniCo under increasing incompleteness, a debiased normalization, and Gaussian jitter, as presented in Fig. 5. All networks are trained for 200 epochs. As incompleteness rises from 25% to 75%, UniCo’s CD increases only from 1.8 to 2.7, whereas strong pointwise baselines double their error to about 6.0. NC for UniCo drops from 0.95 to 0.91, whereas the baselines drop to 0.82. Under a debiased normalization that removes pose and scale canonicalization [56], UniCo degrades to CD 3.9 with NC 0.88, whereas the baselines deteriorate to CD above 14 with NC below 0.65, indicating that UniCo carries substantially less pose and scale bias. With Gaussian noise up to 3%, UniCo remains stable. At a heavy 5% noise level, CD and NC degrade to 3.2 and 0.88, respectively. Fig. 6 further shows view-to-view consistency, where different partial inputs yield a consistent primitive set that supports reliable assembly.
@@
To better understand what drives the strong assembly performance, we further evaluate primitive quality following the established protocol [16, 36, 30]. For all competing methods, primitives are extracted with HPNet [64] for a feasible and fair comparison. Predicted primitives are then matched to ground truth via Hungarian matching. As shown in Tab. 3, we report F-Score@1% (F1), type accuracy (Type), axis difference (Axis), residual error (Res), and coverage (Cov). Axis measures the plane normal or the symmetry axis of cylinders and cones. Res evaluates the fit on 512 points sampled from the ground truth primitive. Cov is the fraction of these points that lie within a distance of 0.01 from the predicted primitive. UniCo attains the best scores on all primitive metrics, consistent with its stronger reconstruction performance.
@@
GRNet [62] 0.215 29.65 25.97 7.12 24.41
PoinTr [69] 0.509 64.41 18.24 2.86 61.60
AdaPoinTr [70] 0.643 79.79 11.34 1.48 78.99
ODGNet [3] 0.659 75.52 12.24 1.78 75.85
SymmComplete [63] 0.629 78.48 12.72 1.54 77.60
@@
Method Year PolyFit [38] KSR [1] COMPOD [45]
@@
GRNet [62] 2020 11.98 19.84 0.769 29.61 9.18 22.01 0.822 10.82 14.19 25.72 0.738 13.15
PoinTr [69] 2021 10.57 16.43 0.822 25.92 8.14 16.33 0.780 30.90 7.82 16.44 0.774 31.34
AdaPoinTr [70] 2023 3.16 7.36 0.920 5.89 3.24 8.86 0.927 0.27 3.25 8.84 0.921 1.32
ODGNet [3] 2024 2.73 6.41 0.933 4.28 2.90 8.56 0.934 0.36 3.22 8.01 0.927 1.05
SymmComplete [63] 2025 3.21 9.15 0.920 4.27 4.23 14.21 0.907 0.38 4.56 14.02 0.895 4.14
PaCo [7] 2025 1.87 4.09 0.943 0.48 1.91 4.14 0.940 0.25 1.94 4.42 0.940 0.25